OpenStudy (anonymous):

Find the volume of the frustum of a regular pyramid with hexagonal bases whose sides are 4 cm and 6 cm and a height of 12 cm

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@kx2bay

OpenStudy (anonymous):

The volume of a pyramid is (1/3) × (area of the base) × height,now find the area

OpenStudy (anonymous):

this is a frustum, not a full pyramid 

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

pffft i know am telling her stepwise The volume of the frustum is then the difference of the volumes of two hexagonal pyramids

OpenStudy (anonymous):

Area of the Base (larger) hexagon = \[\frac{ 3\sqrt{3} }{ 2 } a ^{2}\] and a = 6 can you work it out? what did you get for it?

OpenStudy (anonymous):

93.53

OpenStudy (anonymous):

then what next

OpenStudy (anonymous):

excellent! Can you find the area of the smaller base where a = 4 ?

OpenStudy (anonymous):

41.57 then what next step

OpenStudy (anonymous):

good. http://www.ditutor.com/solid_gometry/frustum_pyramid.html Volume of the frustum = \[V = \frac{ h }{ 3 } ( A + A' + \sqrt{A.A'} )\] h = 12 A = 95.53 A' = 41.57 V = ?

OpenStudy (anonymous):

sorry typo A = 93.53

OpenStudy (anonymous):

v = 789.8

OpenStudy (anonymous):

whats the unit

OpenStudy (anonymous):

cm squared well done

OpenStudy (anonymous):

oh okay thanks godbless always how about this A pyramid whose base is enclosed by a regular hexagon of side 5 cm and whose altitude is 25 m is cut by a plane parallel to the base and 5 cm from the base what is the volume of the frustum formed?

OpenStudy (anonymous):

post the second question seperately, not here

OpenStudy (anonymous):

okay